no. of molecules and atoms in 54gm of H2O

Grade:12

9 Answers

Avinash
8 years ago
No. of moles = wt/GMwt =54/18 = 3

No. of atoms or moleules = 3 x Avagadro no. =3 mol x 6.022 x 1023mol-1 =18.066 x 1023 mol-1
vindhya
8 years ago
will the no. of atoms and no. of molecules be same?
please tell me Avinash sir
Avinash
8 years ago

Yes,no. of atoms and no.of molecules be same since it contains avagadro no of molecules or avagadro no of atoms that mutliplied with no. of moles. so the count will be equal
Prajwal Kavad
8 years ago
no AVINASH sir is wrong , no of molecules are 3xNa and no, of atoms are 9xNa sir there seems to be some problem with you basics
vindhya
8 years ago
avinash sir, please tell me, is prajwal kavad correct?
vindhya
8 years ago
please tell me prajwal kavad, why is it that no. of atoms are 9 x Na ?
kindly explain your answer, if you don't mind.
Avinash
8 years ago
Yes prajwal kavad is right. I’m giving answer for tri atoms and the giving the second reply instead in the thought process of calculation for mono atom.

Vindhya, jst multiply (2 hydrogen atom + 1 oxygen atom) x total no. of molecules

No. of moles = wt/GMwt =54/18 = 3

No. of moleules = 3 x Avagadro no. =3 mol x 6.022 x 1023mol-1=18.066 x 1023

No. of atoms = 3 x18.066 x 1023 = 54.198 x1023
vindhya
8 years ago
thanks a lot to prajwal kavad  for correction and of course Avinash sir  
for a good explanation.
Prajwal Kavad
8 years ago
my pleasure